OpenStudy (anonymous):

Find the area of the base of a pyramid if V = 990 and h = 210.

OpenStudy (anonymous):

volume = 1/3 * h * area base 990 = 1/3 * 210 * a solve for a and you've got it

OpenStudy (anonymous):

can u continue from here?

OpenStudy (anonymous):

Yes Thank You So Much.

OpenStudy (anonymous):

yw